# Conch: Competitive Debate Analysis via Visualizing Clash Points and Hierarchical Strategies

### v1

@@ 6.1 Study Design
Data and tasks. The video and text data from the ICDI were used as the experimental material, as detailed in Table 1. This competition was chosen because it contains richer clash points and disagreements. Moreover, since the data is in Chinese, it reduces participants’ cross-language comprehension burden. We designed seven debate analysis tasks (Table 2) to measure the completion rate of debate analysis from three aspects: what to debate (T1-T3), how to debate (T4-T5) and comprehensive analysis (T6-T7). Following these tasks, we assessed participants’ cognitive load through a ten-item questionnaire for the measurement of IL (Intrinsic Load), EL (Extraneous Load),and GL (Germane Load) [36]. The detailed content of this questionnaire can be found in the supplementary materials. Additionally, a Likert scale (1 = strongly disagree, 7 = strongly agree) was designed for participants in the experimental group to evaluate the effectiveness, visual design, interaction and usability of Conch (Q1 - Q12 in Fig. 8).
@@ C. Comprehensive Analysis
Participants. We recruited 27 debaters and debate coaches to participate in the user study (7 female, 20 male; age: $\text{Mean}(M)=21.0$, Standard Deviation($\text{Standard Deviation}(SD)=2.4$). All of them have 1-9 years of related experience in debate and debate analysis. Specifically, 19 of them have 1-3 years’ experience, 4 have 3-5 years’ experience and 4 have more than 5 years’ experience. Among these participants, two (E3, E4) are world champions in Chinese debate competitions. Based on their length of debate experience, we randomly divided them into three groups through stratified sampling, ensuring similar experience levels within each group. Additionally, another expert (E5), who is also a world champion and the “Best Debater” in Chinese debate competitions, provided feedback on our system.
@@ 6.2 Results
Completion time. As shown in Fig. 10, participants in the experimental group (A) completed the tasks faster than those in the video control group (B) and the text control group (C). Specifically, participants using Conch in the experimental group ($M_{A}=51.1$, $SD_{A}=20.2$) spent 32.8% less time than participants watching debate video ($M_{B}=76.1$, $SD_{B}=17.8$) and 45.9% less than those reading debate text ($M_{C}=94.4$, $SD_{C}=23.8$), which demonstrated statistically significant differences ($p_{B}<0.01$, $p_{C}<0.01$) suggested by the Mann-Whitney U test [38]. Since debate analysis requires a global consideration of the entire content, single task is difficult to be completed and timed independently.
@@
As shown in Fig. 9, we found that IL scores were similar across all three groups (experimental: $3.67\pm 2.17$, control-video: $3.26\pm 1.73$, control-text: $3.96\pm 2.02$), indicating comparable task difficulty and participant knowledge. Using the Mann-Whitney U test, the experimental group had significantly lower EL than the text control group ($1.15\pm 1.12$ vs. $3.15\pm 2.33$, $U=17.5$, $p=0.046$). Compared to the video control group, the experimental group’s EL was lower on average, but the difference was not statistically significant ($1.15\pm 1.12$ vs. $1.59\pm 2.34$, $U=44.5$, $p=0.755$). Additionally, the experimental group’s GL was significantly higher compared to both the video-based group ($7.03\pm 1.85$ vs. $4.22\pm 2.58$, $U=66.5$, $p=0.024$) and the text-based group ($7.03\pm 1.85$ vs. $3.94\pm 2.58$, $U=64.0$, $p=0.042$).
Questionnaire. Fig. 8 reports the results of the user interview questionnaire across four dimensions: effectiveness, visual design, user interaction, and usability. Overall, participants in the experimental group gave high ratings to Conch, with most of the closed-ended questions receiving positive feedback. Specifically, participants expressed satisfaction with the effectiveness and usability of the system, as they rated questions in these dimensions around 6 out of 7. They also provided generally positive ratings for visual design and user interaction, although scores in these two dimensions were slightly less consistent. However, scores for Q5 (“easy to understand”) and Q9 (“interaction is easy for users with any level of debate experience”) were relatively lower. For Q5, participants indicated that the visual design was somewhat challenging to understand due to the introduction of multiple visual components and novel visualization methods. Regarding Q9, participants indicated that the system offered different benefits based on user experience. For beginners, Conch served as a guide to understand the core elements of a debate, such as clash points and strategies; for experts, it functioned as a tool to make their analysis more efficient.
@@ 7 Discussion
Transferability of structured text analysis. We used the competitive debate to characterize the problem domain. The workflow can also be applied to legal text interpretation [55], as shown in Fig. 1. This is because structured texts are characterized by logical frameworks, argumentation, and organized reasoning. We also captured the strategies and patterns used by speakers during argumentation. This enriched our understanding of content interaction and further supported the transferability of our approach.

### v2

@@ 6.1 Study Design
Data and tasks. The video and text data from the ICDI were used as the experimental material, as detailed in Table˜1. This competition was chosen because it contains richer clash points and disagreements. Moreover, since the data is in Chinese, it reduces participants’ cross-language comprehension burden. We designed seven debate analysis tasks (Table˜2) to measure the completion rate of debate analysis from three aspects: what to debate (T1-T3), how to debate (T4-T5) and comprehensive analysis (T6-T7). Following these tasks, we assessed participants’ cognitive load through a ten-item questionnaire for the measurement of IL (Intrinsic Load), EL (Extraneous Load),and GL (Germane Load) [36]. The detailed content of this questionnaire can be found in the supplementary materials. Additionally, a Likert scale (1 = strongly disagree, 7 = strongly agree) was designed for participants in the experimental group to evaluate the effectiveness, visual design, interaction and usability of Conch (Q1 - Q12 in Fig.˜8).
@@ C. Comprehensive Analysis
Participants. We recruited 27 debaters and debate coaches to participate in the user study (7 female, 20 male; age: $\text{Mean}(M)=21.0$, $\text{Standard Deviation}(SD)=2.4$). All of them have 1-9 years of related experience in debate and debate analysis. Specifically, 19 of them have 1-3 years’ experience, 4 have 3-5 years’ experience and 4 have more than 5 years’ experience. Among these participants, two (E3, E4) are world champions in Chinese debate competitions. Based on their length of debate experience, we randomly divided them into three groups through stratified sampling, ensuring similar experience levels within each group. Additionally, another expert (E5), who is also a world champion and the “Best Debater” in Chinese debate competitions, provided feedback on our system.
@@ 6.2 Results
Completion time. As shown in Fig.˜10, participants in the experimental group (A) completed the tasks faster than those in the video control group (B) and the text control group (C). Specifically, participants using Conch in the experimental group ($M_{A}=51.1$, $SD_{A}=20.2$) spent 32.8% less time than participants watching debate video ($M_{B}=76.1$, $SD_{B}=17.8$) and 45.9% less than those reading debate text ($M_{C}=94.4$, $SD_{C}=23.8$), which demonstrated statistically significant differences ($p_{B}<0.01$, $p_{C}<0.01$) suggested by the Mann-Whitney U test [38]. Since debate analysis requires a global consideration of the entire content, single task is difficult to be completed and timed independently.
@@
As shown in Fig.˜9, we found that IL scores were similar across all three groups (experimental: $3.67\pm 2.17$, control-video: $3.26\pm 1.73$, control-text: $3.96\pm 2.02$), indicating comparable task difficulty and participant knowledge. Using the Mann-Whitney U test, the experimental group had significantly lower EL than the text control group ($1.15\pm 1.12$ vs. $3.15\pm 2.33$, $U=17.5$, $p=0.046$). Compared to the video control group, the experimental group’s EL was lower on average, but the difference was not statistically significant ($1.15\pm 1.12$ vs. $1.59\pm 2.34$, $U=44.5$, $p=0.755$). Additionally, the experimental group’s GL was significantly higher compared to both the video-based group ($7.03\pm 1.85$ vs. $4.22\pm 2.58$, $U=66.5$, $p=0.024$) and the text-based group ($7.03\pm 1.85$ vs. $3.94\pm 2.58$, $U=64.0$, $p=0.042$).
Questionnaire. Fig.˜8 reports the results of the user interview questionnaire across four dimensions: effectiveness, visual design, user interaction, and usability. Overall, participants in the experimental group gave high ratings to Conch, with most of the closed-ended questions receiving positive feedback. Specifically, participants expressed satisfaction with the effectiveness and usability of the system, as they rated questions in these dimensions around 6 out of 7. They also provided generally positive ratings for visual design and user interaction, although scores in these two dimensions were slightly less consistent. However, scores for Q5 (“easy to understand”) and Q9 (“interaction is easy for users with any level of debate experience”) were relatively lower. For Q5, participants indicated that the visual design was somewhat challenging to understand due to the introduction of multiple visual components and novel visualization methods. Regarding Q9, participants indicated that the system offered different benefits based on user experience. For beginners, Conch served as a guide to understand the core elements of a debate, such as clash points and strategies; for experts, it functioned as a tool to make their analysis more efficient.
@@ 7 Discussion
Transferability of structured text analysis. We used the competitive debate to characterize the problem domain. The workflow can also be applied to legal text interpretation [55], as shown in Fig.˜1. This is because structured texts are characterized by logical frameworks, argumentation, and organized reasoning. We also captured the strategies and patterns used by speakers during argumentation. This enriched our understanding of content interaction and further supported the transferability of our approach.